OpenStudy (anonymous):

what are factors of 20?

OpenStudy (tkhunny):

\[\sqrt{20}\cdot\sqrt{20}\] Prime Factors? Integers? Whole Numbers?

OpenStudy (anonymous):

N is the set of real numbers that are factors of 20.

OpenStudy (tkhunny):

Generally, Prime Factors are needed. Look for anything that divides 20. '2' comes to mind. 20 2*10 2 also divides 10 2*2*5 5 is a prime number, and we're done.